A company produces reuseable water bottles. Their technology allows one worker to produce 10 water bottles in an hour. Hiring a second worker would allow them to produce 16 total water bottles in an hour and three workers can produce 18 total water bottles. What is the marginal benefit of hiring the second worker? A 10 B 16 C 6 D 2
Answer #1

Answer

marginal product of n th worker =total output of n workers - total output of n-1 workers
MP(n)=TP(n)-TP(n-1)

MP(2)=16-10=6 units

the marginal benefit of hiring the second worker is 6 units

Option C